OpenStudy (anonymous):

can someone show me the steps to solving this equation 5/13 t = -9

OpenStudy (anonymous):

Multiply each side by 13t. So you get 5=-9 * 13t Now divide each side by -9. -5/9= 13t. Final step is to divide by 13. -5/117=t